HIGH VIBRATIONAL LEVELS OF O_{2}(b^{1}Σ_{g}^{+}) AND O_{2}(a^{1}Δ_{g})

K. S. KALOGERAKIS, \underline{D. L. HUESTIS}, P. C. COSBY, T. G. SLANGER, R. A. COPELAND , Molecular Physics Laboratory, SRI International, Menlo Park, CA 94025; A. TOTTH, Bowdoin College, Brunswick, Me.

Relaxation of laser-excited O_{2}(A^{3}Σ_{u}^{+}, v = 6 − 10) in collisions with O_{2} populates high vibrational levels of O_{2}(b^{1}Σ_{g}^{+}) and O_{2}(a^{1}Δ_{g}). Previous work on the spectroscopy of O_{2} emissions from the Earth’s night atmosphere has enabled us to assign the Q-branches of thirteen (v’,v”) bands in the (2+1) REMPI spectra of O_{2}(b^{1}Σ_{g}^{+}, v” = 10 − 15) and O_{2}(a^{1}Δ_{g}, v” = 16 − 19), through the intermediate Rydberg levels 3dπ ^{1}Σ_{g}^{+}(v′ = 2 − 5) and ^{1}Δ_{g}(v′ = 4, 5), repectively.

The derived spectroscopic constants for O_{2}(b^{1}Σ_{g}^{+}, v” = 10 − 13) agree with our previous work. Improved values are obtained for v” = 14, 15. Our constants for 3dπ ^{1}Σ_{g}^{+}(v′ = 5) are the first available. Our spectra are better calibrated and our constants for the v′ = 2 − 4 levels significantly more precise than the 1988-1992 work in the groups of Chupka and Houston.

The energies of O_{2}(a^{1}Δ_{g}, v” = 16 − 19) were known previously only to within about 80 cm^{−1} from the 1995 electron scattering work of Allan. The highest level known accurately from nightglow spectra is v” = 11. Our spectral analysis gives precise rotational constants and ΔG values for these vibrational levels and for 3dπ ^{1}Δ_{g}(v′ = 4, 5). The absolute term energies are uncertain by about 5 cm^{−1} based on earlier work in the groups of Chupka and Houston.

Supported by the NASA Geospace Sciences and Planetary Atmospheres programs and by the NSF Aeronomy and Physics/REU programs.